# A rotational hyperbolic theory for surface homeomorphisms

### v1

@@
Abstract
We develop a rotational hyperbolic theory for surface homeomorphisms. We use the equivalence relation on ergodic measures that have nontrivial rotational behaviour defined in [GSGL24] to define a rotational counterpart of homoclinic classes. These allows to produce a network of horseshoes representing the whole rotational behaviour f the homeomorphism. We also study the counterpart of heteroclinic connections and give 5 different characterizations of such connections.
@@ Abstract
1 Introduction
@@
$\rho(\mu)=\int_{S}[a_{y}]\,\mathrm{d}\mu(y)=\lim\limits_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}[a_{f^{i}(x)}].$ (2)
@@
$\lim\limits_{n\to+\infty}\frac{1}{n}d\big(\widetilde{f}^{n}(\widetilde{z}),\widetilde{\gamma}_{\widetilde{z}}(n\vartheta_{\mu})\big)=\lim\limits_{n\to+\infty}\frac{1}{n}d\big(\widetilde{f}^{-n}(\widetilde{z}),\widetilde{\gamma}_{\widetilde{z}}(-n\vartheta_{\mu})\big)=0.$ (3)
@@
$\rho_{i}=\{\rho(\mu)\mid\mu\in\mathcal{N}_{i}\}\qquad\text{and}\qquad\dot{\Lambda}_{i}=\bigcup_{\mu\in\mathcal{N}_{i}}\dot{\Lambda}_{\mu}.$ (4)
@@
Figure 1: Idea of the statement of Theorem 3.11: if there is a trajectory under the isotopy like the one in the left of the figure in the space of leaves, then there exists two rotational horseshoes for $f$ having a heteroclinic connection.
@@ 2.1 Forcing theory
Figure 2: Example of $^\widehat{\mathcal{F}}$-transverse intersection.
@@
2.2 Classification of ergodic rotation sets
@@
2.3 Bounded deviations in homotopy
@@
3 Heteroclinic horseshoes in forcing theory
3.1 Markovian intersections
@@
Figure 3: A pre-Markovian intersection (left) and a Markovian intersection(right). The horizontal sub-rectangle for the pre-markovian intersection is denoted $R^{\prime}_{1}$.
@@
3.2 Heteroclinic connections of horseshoes and rotation sets
@@
$\frac{p_{m}^{q}k_{m}r_{i_{m}}}{q}\underset{q\to+\infty}{\longrightarrow}\sigma_{i_{m}}.$ (6)
@@
$\widetilde{f}^{\,\sum_{j=0}^{k}n(w_{j})}(\widetilde{x})\in T(w_{0})T(w_{1})\cdots T(w_{k})\,R_{e(w_{k})}.$ (8)
@@
3.3 Creation of heteroclinic connections of horseshoes by forcing theory
@@
Figure 4: Beginning of the proof of Theorem 3.11 for $k_{1}=3$: construction of the rectangle $\widehat{R}_{1}$ (top) and Markovian intersections of the image $\widehat{f}^{3}(\widehat{R}_{1})$ with $T_{1}\widehat{R}_{1}$, $T_{1}^{2}\widehat{R}_{1}$ and $T_{1}^{3}\widehat{R}_{1}$ (bottom).
@@
Figure 5: Proof of Theorem 3.11 in the case $k_{1}=k_{2}=2$: construction of the rectangles $\widehat{R}_{1}$ and $\widehat{R}_{2}$.
@@
Figure 6: Proof of Theorem 3.11: the Markovian intersection $\widehat{f}^{n}(\widehat{R}_{1})\cap\widehat{R}_{2}$.
@@
4 Heteroclinic connections between chaotic classes
4.1 A graph $G$
@@
$I^{t_{\omega}^{-}}_{\mathcal{F}}(z^{\prime}_{\omega})=I^{t_{\omega}^{+}+nq^{\prime}_{\omega}}_{\mathcal{F}}(z^{\prime}_{\omega}),$ (9)
@@
$\widetilde{f}^{q_{\omega}}(R_{\omega})\cap_{M}T_{\omega}R_{\omega}\qquad\text{and}\qquad\frac{[T_{\omega}]}{q_{\omega}}=r_{\omega}$ (10)
@@
4.2 Connections between chaotic classes
@@
$\frac{1}{n^{1,+}_{k}-n^{1,-}_{k}}\sum_{i=n^{1,-}_{k}}^{n^{1,+}_{k}-1}\delta_{f^{i}(x_{k})}\xrightharpoonup[k\to+\infty]{}\mu_{1}\quad\textrm{and}\quad\frac{1}{n^{2,+}_{k}-n^{2,-}_{k}}\sum_{i=n^{2,-}_{k}}^{n^{2,+}_{k}-1}\delta_{f^{i}(x_{k})}\xrightharpoonup[k\to+\infty]{}\mu_{2}$ (11)
@@
$d\big(\widetilde{f}^{-n}(\widetilde{x}),\widetilde{\gamma}_{\omega}(-n\vartheta_{i})\big)\leq D\qquad\text{and}\qquad d\big(\widetilde{f}^{n}(\widetilde{x}),\widetilde{\gamma}_{\omega^{\prime}}(n\vartheta_{j})\big)\leq D.$ (12)
@@
$d\big(\widetilde{f}^{-n}(\widetilde{x}_{i}),\widetilde{\gamma}_{\omega}(-n\vartheta_{i})\big)\leq D\qquad\text{and}\qquad d\big(\widetilde{f}^{n}(\widetilde{x}_{j}),\widetilde{\gamma}_{\omega^{\prime}}(n\vartheta_{j})\big)\leq D.$ (13)
@@ 4.3 Further properties of $\to$
4.4 A graph $\mathcal{T}$ associated to the surface
@@
Figure 7: An example of the classes $(\mathcal{N}_{i})_{i\in I_{\mathrm{h}}\cup I^{1}}$ (defined in Definitions 1.6 and 1.7) of a homeomorphism of a closed genus 6 surface. The classes $(\mathcal{N}_{i})_{1\leq i\leq 5}$ belong to $I_{\mathrm{h}}$, while the classes $\mathcal{N}_{6}$ and $\mathcal{N}_{7}$ belong to $I^{1}$. The geodesic lamination $\dot{\Lambda}_{6}$ is made of a single closed geodesic, while $\dot{\Lambda}_{7}$ is a minimal geodesic lamination with non closed leaves ([GSGL24, Theorem D].
@@
4.5 Some open invariant sets associated to the graph $\mathcal{T}$

### v2

@@
Abstract.
We develop a rotational hyperbolic theory for surface homeomorphisms. We use the equivalence relation on ergodic measures that have nontrivial rotational behaviour defined in [GSGL24] to define a rotational counterpart of homoclinic classes. These allows to produce a network of horseshoes representing the whole rotational behaviour of the homeomorphism. We also study the counterpart of heteroclinic connections and give 5 different characterizations of such connections.
@@ Abstract.
P.-A. G. thanks the Jean-Christophe Yoccoz international laboratory CNRS/IMPA for the semester in Brazil during which the ideas of this work were born.
1. Introduction
@@
(2) $\rho(\mu)=\int_{S}[a_{y}]\,\mathrm{d}\mu(y)=\lim\limits_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}[a_{f^{i}(x)}].$
@@
(3) $\lim\limits_{n\to+\infty}\frac{1}{n}d\big(\widetilde{f}^{n}(\widetilde{z}),\widetilde{\gamma}_{\widetilde{z}}(n\vartheta_{\mu})\big)=\lim\limits_{n\to+\infty}\frac{1}{n}d\big(\widetilde{f}^{-n}(\widetilde{z}),\widetilde{\gamma}_{\widetilde{z}}(-n\vartheta_{\mu})\big)=0.$
@@
(4) $\rho_{i}=\{\rho(\mu)\mid\mu\in\mathcal{N}_{i}\}\qquad\text{and}\qquad\dot{\Lambda}_{i}=\bigcup_{\mu\in\mathcal{N}_{i}}\dot{\Lambda}_{\mu}.$
@@
Figure 1. Idea of the statement of Theorem 3.11: if there is a trajectory under the isotopy like the one in the left of the figure in the space of leaves, then there exists two rotational horseshoes for $f$ having a heteroclinic connection.
@@ 2.1. Forcing theory
Figure 2. Example of $^\widehat{\mathcal{F}}$-transverse intersection.
@@
2.2. Classification of ergodic rotation sets
@@
2.3. Bounded deviations in homotopy
@@
3. Heteroclinic horseshoes in forcing theory
3.1. Markovian intersections
@@
Figure 3. A pre-Markovian intersection (left) and a Markovian intersection(right). The horizontal sub-rectangle for the pre-markovian intersection is denoted $R^{\prime}_{1}$.
@@
3.2. Heteroclinic connections of horseshoes and rotation sets
@@
(6) $\frac{p_{m}^{q}k_{m}r_{i_{m}}}{q}\underset{q\to+\infty}{\longrightarrow}\sigma_{i_{m}}.$
@@
(8) $\widetilde{f}^{\,\sum_{j=0}^{k}n(w_{j})}(\widetilde{x})\in T(w_{0})T(w_{1})\cdots T(w_{k})\,R_{e(w_{k})}.$
@@
3.3. Creation of heteroclinic connections of horseshoes by forcing theory
@@
Figure 4. Beginning of the proof of Theorem 3.11 for $k_{1}=3$: construction of the rectangle $\widehat{R}_{1}$ (top) and Markovian intersections of the image $\widehat{f}^{3}(\widehat{R}_{1})$ with $T_{1}\widehat{R}_{1}$, $T_{1}^{2}\widehat{R}_{1}$ and $T_{1}^{3}\widehat{R}_{1}$ (bottom).
@@
Figure 5. Proof of Theorem 3.11 in the case $k_{1}=k_{2}=2$: construction of the rectangles $\widehat{R}_{1}$ and $\widehat{R}_{2}$.
@@
Figure 6. Proof of Theorem 3.11: the Markovian intersection $\widehat{f}^{n}(\widehat{R}_{1})\cap\widehat{R}_{2}$.
@@
4. Heteroclinic connections between chaotic classes
4.1. A graph $G$
@@
(9) $I^{t_{\omega}^{-}}_{\mathcal{F}}(z^{\prime}_{\omega})=I^{t_{\omega}^{+}+nq^{\prime}_{\omega}}_{\mathcal{F}}(z^{\prime}_{\omega}),$
@@
(10) $\widetilde{f}^{q_{\omega}}(R_{\omega})\cap_{M}T_{\omega}R_{\omega}\qquad\text{and}\qquad\frac{[T_{\omega}]}{q_{\omega}}=r_{\omega}$
@@
4.2. Connections between chaotic classes
@@
(11) $\frac{1}{n^{1,+}_{k}-n^{1,-}_{k}}\sum_{i=n^{1,-}_{k}}^{n^{1,+}_{k}-1}\delta_{f^{i}(x_{k})}\xrightharpoonup[k\to+\infty]{}\mu_{1}\quad\textrm{and}\quad\frac{1}{n^{2,+}_{k}-n^{2,-}_{k}}\sum_{i=n^{2,-}_{k}}^{n^{2,+}_{k}-1}\delta_{f^{i}(x_{k})}\xrightharpoonup[k\to+\infty]{}\mu_{2}$
@@
(12) $d\big(\widetilde{f}^{-n}(\widetilde{x}),\widetilde{\gamma}_{\omega}(-n\vartheta_{i})\big)\leq D\qquad\text{and}\qquad d\big(\widetilde{f}^{n}(\widetilde{x}),\widetilde{\gamma}_{\omega^{\prime}}(n\vartheta_{j})\big)\leq D.$
@@
(13) $d\big(\widetilde{f}^{-n}(\widetilde{x}_{i}),\widetilde{\gamma}_{\omega}(-n\vartheta_{i})\big)\leq D\qquad\text{and}\qquad d\big(\widetilde{f}^{n}(\widetilde{x}_{j}),\widetilde{\gamma}_{\omega^{\prime}}(n\vartheta_{j})\big)\leq D.$
@@ 4.3. Further properties of $\to$
4.4. A graph $\mathcal{T}$ associated to the surface
@@
Figure 7. An example of the classes $(\mathcal{N}_{i})_{i\in I_{\mathrm{h}}\cup I^{1}}$ (defined in Definitions 1.6 and 1.7) of a homeomorphism of a closed genus 6 surface. The classes $(\mathcal{N}_{i})_{1\leq i\leq 5}$ belong to $I_{\mathrm{h}}$, while the classes $\mathcal{N}_{6}$ and $\mathcal{N}_{7}$ belong to $I^{1}$. The geodesic lamination $\dot{\Lambda}_{6}$ is made of a single closed geodesic, while $\dot{\Lambda}_{7}$ is a minimal geodesic lamination with non closed leaves ([GSGL24, Theorem D].
@@
4.5. Some open invariant sets associated to the graph $\mathcal{T}$